OpenStudy (anonymous):

Solve 2x2 = 14x + 20

OpenStudy (anonymous):

-16/14

OpenStudy (anonymous):

just do what you are asked.. 2*2=4, so 4=14x+20 or: 14x+20=4 -20 from both sides: 14x=-16 divide both sides by 14: x=-16/14 simplify: x=-8/7